# SceneComplete: Open-World 3D Scene Completion in Cluttered Real World Environments for Robot Manipulation

### v4

@@ II Related Work
Figure 2: Overview of the SceneComplete pipeline. Starting from a single RGB-D input, the system produces a set of object meshes registered with the input 3D scan, yielding a complete 3D scene reconstruction. The pipeline consists of six key phases: (1) An RGB image is fed into a vision-language model to enumerate and describe objects, (2) object descriptions and the RGB image are processed by a grounded segmentation model to generate object masks, (3) occluded regions are completed via image inpainting model adapted to output single fully observable objects on a white background, (4) the inpainted 2D images are passed into an image-to-3D model to produce object meshes, (5) object meshes are scaled according to the segmented partial point cloud, and (6) mesh poses are adjusted within the 3D coordinate frame of the original scan using 6DOF pose estimation. Each step leverages pre-trained open world large vision models, enabling scalability and benefiting from future model improvements.
@@ IV-A Scene Reconstruction and Grasping
Collision-free Grasping: We perform a basic test of the utility of SceneComplete on an important downstream task of grasping, by using an antipodal grasp generation method [3] to generate grasps on the objects in the reconstructed scene. To illustrate the importance of whole-scene reconstruction on grasping, we ask the question: of these potential grasps, which ones are in collision with the ground-truth scene? Concretely, for a reconstructed scene, we (a) sample a set $G$ of collision-free grasps using antipodal sampling, (b) evaluate grasps in $G$ that would cause a collision in the ground truth scene to obtain a subset $G^{\prime}$, and compute the Grasp Collision metric GC as ($\lvert G^{\prime}\rvert/\lvert G\rvert$) which is the percentage of grasps that the reconstruction would allow, that in fact collide, similar to the metric adopted by other works [31, 19]. In our case, we set $G$ to 40.
@@ V Discussion
VI Acknowledgments
We gratefully acknowledge support from NSF grant 2214177; from AFOSR grant FA9550-22-1-0249; from ONR MURI grants N00014-22-1-2740 and N00014-24-1-2603; from the MIT Quest for Intelligence; and from the Robotics and Artificial Intelligence Institute.

### v5

@@ Abstract
Figure 1: (a) takes as input a single RGB-D image of a given scene, visualized here as a point cloud; (b) produces high-quality fully completed, accurately segmented object meshes in scenes with substantial occlusion and novel objects; and (c) enables downstream dexterous manipulation that requires accurate complete shape information.
Manuscript received: June 14, 2025; Revised: September 11, 2025; Accepted: October 14, 2025This paper was recommended for publication by Editor Júlia Borràs Sol upon evaluation of the Associate Editor and Reviewers’ comments.We gratefully acknowledge support from NSF grant 2214177; from AFOSR grant FA9550-22-1-0249; from ONR MURI grants N00014-22-1-2740 and N00014-24-1-2603; from the MIT Quest for Intelligence; and from the Robotics and Artificial Intelligence Institute.1Aditya Agarwal, Bipasha Sen, Tomás Lozano-Pérez, and Leslie Pack Kaelbling are with Computer Science and Artificial Intelligence Laboratory, MIT. {adityaag, bise, tlp, lpk}@mit.edu2Gaurav Singh is with the Department of Computer Science, Brown University. gaurav@brown.eduDigital Object Identifier (DOI): see top of this page.
@@ I Introduction
Figure 2: Overview of the SceneComplete pipeline. Starting from a single RGB-D input, the system produces a set of object meshes registered with the input 3D scan, yielding a complete 3D scene reconstruction. The pipeline consists of six key phases: (1) An RGB image is fed into a VLM to enumerate and describe objects, (2) object descriptions and the RGB image are processed by a grounded segmentation model to generate object masks, (3) occluded regions are completed via image inpainting model adapted to output single fully observable objects on a white background, (4) the inpainted 2D images are passed into an image-to-3D model to produce object meshes, (5) object meshes are scaled according to the segmented partial point cloud, and (6) mesh poses are adjusted within the 3D coordinate frame of the original scan using 6DOF pose estimation. Each step leverages pre-trained open world large vision models, enabling scalability and benefiting from future model improvements.
@@ IV-A Scene Reconstruction and Grasping
Collision-free Grasping: We perform a basic test of the utility of SceneComplete on an important downstream task of grasping, by using an antipodal grasp generation method [3] to generate grasps on the objects in the reconstructed scene. To illustrate the importance of whole-scene reconstruction on grasping, we ask the question: of these potential grasps, which ones are in collision with the ground-truth scene? Concretely, for a reconstructed scene, we (a) sample a set $G$ of collision-free grasps using antipodal sampling, (b) evaluate grasps in $G$ that would cause a collision in the ground truth scene to obtain a subset $G^{\prime}$, and compute the Grasp Collision metric GC as ($\lvert G^{\prime}\rvert/\lvert G\rvert$) which is the percentage of grasps that the reconstruction would allow, that in fact collide, similar to the metric adopted by [31, 19]. In our case, $G$ is set to 40.